7. An infusion pump is a medical device that delivers fluids into a patient's body in controlled amounts. The concentration of drug A in the solution is 0.96 mg mL . The pump rate is 12.8 mL hour . If the pump is running for 4 hours 30 minutes how many milligrams of drug A will be delivered ? Helpful info (helps you understand what is happening): 12.8 ?? ℎ??? means that each hour 12.8 mL of the solution gets delivered. 0.96 ?? ?? means that each mL of the solution contains 0.96 mg of the drug; Start with converting 4 hours 30 minutes to a decimal format.
Answer #1

Pump ran for 4 h 30 min.We know that 1 h = 60 min

4 h 30 min = 4 h + 30 min = 4 h + 30 min x (1 h/60 min) = 4 h + 0.5 h = 4.5 h

Pump rate = 12.8 mL/h

Volume of solution delivered = Pump rate x Total Time = 12.8 mL/h x 4.5 h = 57.6 mL

1 mL of solution contains 0.96 mg drug.

Thus, total drug delivered = Volume of solution delivered x 0.96 mg/1 mL = 57.6 mL x (0.96 mg/1 mL) = 55.3 mg  

Thus, the answer will be 55.3 mg which when rounded to 2 significant figure will become 55. mg